OpenStudy (anonymous):

After walking for 10 minutes you are 80% of the way to school. Use a proportion to find out how long it will take you to walk to school, assuming you are walking at a constant rate and make no stops. Ok. Im. Stuck. :(

OpenStudy (anonymous):

80/100=0.8 and thats all i got so far...

OpenStudy (anonymous):

if it takes 10 minute to take 80% of the way then it takes 5 min to take 40% of the way( if you divide by 2), and if you divide again, it takes 2.5 min to take 20%. So if you add these up 80%=10min, 20%=2.5min then it takes 12.5min to take 100%

OpenStudy (anonymous):

Oh ok :) haha i feel stupid /: so 80%/2=40%(10 minutes) then 40/2=20%(5 minutes) then 5mins/2=2.5mins :D Thanks!

OpenStudy (anonymous):

80% of the way is done by walking for 10 minutes. To be at school, you must complete your walking, and that's 100% progress, and means that you must walk farther(or longer). The answer is 100/80 x 10 minutes = 12,5 minutes